Monty Company acquired a plant asset at the beginning of Year 1. The asset has an estimated service life of 5 years. An employee has prepared depreciation schedules for this asset using three different methods to compare the results of using one method with the results of using other methods. You are to assume that the following schedules have been correctly prepared for this asset using (1) the straight-line method, (2) the sum-of-the-years'-digits method, and (3) the double-declining-balance method.

Year

Straight-Line

Sum-of-the-
Years'-Digits

Double-Declining-
Balance

1 $10,620 $17,700 $23,600
2 10,620 14,160 14,160
3 10,620 10,620 8,496
4 10,620 7,080 5,098
5 10,620 3,540 1,746
Total $53,100 $53,100 $53,100


Answer the following questions.

What is the cost of the asset being depreciated?

Cost of asset

What amount, if any, was used in the depreciation calculations for the salvage value for this asset?

Salvage value
Answer #1
Double-Declining-rate 40% =1/5*2
a
Cost of asset 59000 =23600/40%
b
Salvage value 5900 =59000-53100